# Numerical simulation of type II primordial black hole formation

### 12 Jan 2024 v1

@@ 1 Introduction
Kopp, Hofmann, and Weller (KHW) [1] investigated PBH formation associated with the type II fluctuations by employing the Lemaître–Tolman–Bondi (LTB) solution, which is the analytic solution for the dust fluid ($w=0$). They investigated the causal structure of the spacetime generated from a specific profile of the type II fluctuation. They explicitly showed that it describes a spacetime of BH formation, which they call ‘type II PBH.’ In addition, they revealed distinct properties of the type II PBH from ‘type I PBH,’ which arises from typical type I fluctuations with the amplitude over the threshold. We aim to investigate the dynamics and the spacetime structure generated from type II fluctuations in the radiation-dominated universe, namely, to clarify the impact of the existence of the radiation pressure.
@@ 3.3 Spacetime structures of type I/II PBH
The crucial difference between the right and left panels in Fig. 9 is the existence of the bifurcating trapping horizons $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(0;*)$, which is described by the intersection points of the two trapping horizon trajectories in Fig. 8. In this paper, we propose to use this difference to classify PBH formation processes into types I and II; that is, we identify the PBH as of type I if there is no bifurcating trapping horizon and of type II if there appears at least one bifurcating trapping horizon during the time evolution associated with the intersection point. This crucial difference has already been pointed out by KHW in Ref. [1] for the case of the dust fluid. Unlike the current proposal, the authors in Ref. [1] did not clearly distinguish the definitions of type I/II PBHs and type I/II fluctuations. However, in our simulation with $w=1/3$, we find that the horizon configuration is of type I for the $\mu=1.6$ case, but the initial fluctuation is of type II nevertheless. Therefore, there is a gap between the definitions of type I/II PBHs and type I/II fluctuations. This fact is our motivation for the new definition of the PBH.
@@ 4.1 PBH mass with the Gaussian initial fluctuation
Let us consider how the mass of the PBH depends on the initial amplitude $\mu$ of the fluctuation in the type II regime by using the three-zone model with dust as seen in Fig. 18. The PBH mass is then given by the mass parameter for the Schwarzschild spacetime in the middle patch. If we fix the scale of the spatial curvature of the closed FLRW region, the horizon area of the Schwarzschild region is fixed by the horizon area on the matching surface between the closed FLRW and the Schwarzschild regions. Increasing the initial amplitude, we expect that the closed FLRW region will be extended and the boundary will be shifted more to the right. In the closed FLRW side, the shift of the boundary to the right results in the decreasing area of the horizons. Therefore, the mass parameter for the Schwarzschild region is also expected to decrease. From the above consideration of the model with dust, we may naively guess that the mass of the resultant PBH would decrease with the initial amplitude value for the type II regime. See also Ref. [41]. However, as will be shown soon, this guess is not necessarily true for the radiation case, and the behavior also strongly depends on the profile of the initial fluctuation.
@@
where $\rho_{b}=\rho_{b,0}(t_{0}/t)^{2}$ is the background energy density, and $F$ is a constant that corresponds to the efficiency of accretion, which is commonly numerically found to be of order $\mathcal{O}(1)$. It is important to note that Eq. (4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion. But it will correctly describe the accretion into the PBHs for sufficiently large times once we have a quasi-stationary flow into the PBH. The integrated solution is,
@@ 4.2 PBH mass with the initial fluctuation motivated by the three-zone model
where $1/k$ is the comoving radius of $S^{3}$ associated with the closed FLRW and $0<\chi_{a}\leq\pi/k$ describes the radial coordinate of the edge of the closed FLRW region. Although this expression is relevant for the region $r\leq r_{a}$, we simply extend this expression in the region $r>r_{a}$, and to realize the condition $\zeta(L)=0$, we use the following specific functional form of $\zeta(r)$:
@@
The behavior of the horizon configuration is similar to the case of the Gaussian profile, and we do not explicitly show it. In contrast to the Gaussian profile, the initial PBH mass formed by this profile decreases as the amplitude of fluctuation $\tilde{\mu}$ increases for $\tilde{\mu}>0.65$ as shown in Fig. 15, even though they are classified as type I fluctuations and formations for $\tilde{\mu}\leq 0.88$. This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases with the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of nonlinear perturbation. We also find the gap between the parameter region of type I/II fluctuations and type I/II PBHs, as indicated in Fig. 15.
@@ Appendix A Conformal diagrams of the three-zone model for the dust case
For the underdense region with a perfect fluid with $w=1/3$, it would not be tractable to obtain any analytical and exact solutions of such dynamical inhomogeneous spacetimes with nonlinear sound waves. Therefore, we stick to the three-zone model with dust. Furthermore, for simplicity, we focus on the model where the underdense region is described by the Schwarzschild solution, where the matching is possible without any singular hypersurface for the dust case. In spite of this great simplification, the model provides us with a very powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons. Note that one can also approach the same problem by using the Lemaître-Tolman-Bondi exact dust solution (c.f. [1]).

### 02 Feb 2024 v2

@@ 1 Introduction
Kopp, Hofmann, and Weller (KHW) [1] investigated PBH formation associated with the type II fluctuations by employing the Lemaître–Tolman–Bondi (LTB) solution, which is the analytic solution for the dust fluid ($w=0$). They investigated the causal structure of the spacetime generated from a specific profile of the type II fluctuation. They explicitly showed that it describes a spacetime of BH formation, which they call ‘type II PBH.’ In addition, they revealed distinct properties of the type II PBH from ‘type I PBH,’ which arises from typical type I fluctuations with the amplitude over the threshold. Carr and Harada (2015) [22] analytically discussed type II PBHs with more general equations of state using simplified models. Type I/II fluctuations are also important because the threshold of these two cases gives a possible maximum value of the averaged density perturbation at horizon entry in the simplified models [1, 22]. We aim to investigate the dynamics and the spacetime structure generated from type II fluctuations in the radiation-dominated universe, namely, to clarify the impact of the existence of the radiation pressure.
@@ 3.3 Spacetime structures of type I/II PBH
The crucial difference between the right and left panels in Fig. 9 is the existence of the bifurcating trapping horizons $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(0;*)$, which is described by the intersection points of the two trapping horizon trajectories in Fig. 8. In this paper, we propose to use this difference to classify PBH formation processes into types I and II; that is, we identify the PBH as of type I if there is no bifurcating trapping horizon and of type II if there appears at least one bifurcating trapping horizon during the time evolution associated with the intersection point. This difference has already been pointed out by KHW in Ref. [1] for the case of the dust fluid. Unlike the current proposal, the authors in Ref. [1] did not clearly distinguish the definitions of type I/II PBHs and type I/II fluctuations. However, in our simulation with $w=1/3$, we find that the horizon configuration is of type I for the $\mu=1.6$ case, but the initial fluctuation is of type II nevertheless. Therefore, there is a gap between the definitions of type I/II PBHs and type I/II fluctuations. This fact is our motivation for the new definition of the PBH.
@@ 4.1 PBH mass with the Gaussian initial fluctuation
[…]
@@
where $\rho_{b}=\rho_{b,0}(t_{0}/t)^{2}$ is the background energy density, and $F$ is a constant that corresponds to the efficiency of accretion, which is commonly numerically found to be of order $\mathcal{O}(1)$. It is important to note that Eq. (4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion as Carr and Hawking (1974) [4] already pointed out. But it will correctly describe the accretion into the PBHs for sufficiently large times once we have a quasi-stationary flow into the PBH. The integrated solution is,
@@ 4.2 PBH mass with the initial fluctuation motivated by the three-zone model
where $k^{-1}$ is the comoving radius of $S^{3}$ associated with the closed FLRW and $0<\chi_{a}\leq\pi/k$ describes the radial coordinate of the edge of the closed FLRW region. Although this expression is relevant for the region $r\leq r_{a}$, we simply extend this expression in the region $r>r_{a}$, and to realize the condition $\zeta(L)=0$, we use the following specific functional form of $\zeta(r)$:
@@
The behavior of the horizon configuration is similar to the case of the Gaussian profile, and we do not explicitly show it. In contrast to the Gaussian profile, the initial PBH mass formed by this profile decreases as the amplitude of fluctuation $\tilde{\mu}$ increases for $\tilde{\mu}>0.65$ as shown in Fig. 15, even though they are classified as type I fluctuations and formations for $\tilde{\mu}\leq 0.88$. This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases with the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of non-linear perturbation. We also find the gap between the parameter region of type I/II fluctuations and type I/II PBHs, as indicated in Fig. 15.
@@ Appendix A Conformal diagrams of the three-zone model for the dust case
For the underdense region with a perfect fluid with $w=1/3$, it would not be tractable to obtain any analytical and exact solutions of such dynamical inhomogeneous spacetimes with non-linear sound waves. Therefore, we stick to the three-zone model with dust. Furthermore, for simplicity, we focus on the model where the underdense region is described by the Schwarzschild solution, where the matching is possible without any singular hypersurface for the dust case. In spite of this great simplification, the model provides us with a very powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons. Note that one can also approach the same problem by using the Lemaître-Tolman-Bondi exact dust solution (c.f. [1]).